OpenStudy (anonymous):

show that x^2+ax+y^2+by+c=0 is the equation of a circle if and only if a^2+b^2-4c>0

OpenStudy (dumbcow):

use completing the square to put equation of circle in standard form (x+a/2)^2 + (y+b/2)^2 = (a^2/4)+(b^2/4)-c r^2 = (a^2/4)+(b^2/4)-c > 0 multiply equation by 4 -> a^2 +b^2 -4c > 0

OpenStudy (anonymous):

ok so am curious, why must it be greater than 0 ?

OpenStudy (dumbcow):

the square of a number is always positive thus greater than 0. also a circle must have a positive radius

OpenStudy (anonymous):

ok cool ty for the help